a ray of light is incident at an angle of 40 degree on a prism of refracting angle 60 degrees is deviated through an angle of 38 degrees calculate angle of emergence

Grade:11

1 Answers

aishwarya jamdar
6 years ago
By prisam formula,
i+e=A+delta
angle of incedence(i)=40degree, refractin angle of prism(A)=60degree, angle of deviation(delta)=38
e=A+delta-i
  =60+38-40=58degree